how do u do these types of equations 16z - 18 = 4 + 5z

Answer by boilpoil(127)

16z-18=4+5z
11z=22 --- Moved and calculated: -18 to RHS, 4+18 = 22. +5z to LHS, 16z-5z = 11z
z=2 --- Divided 11z and 22 by 11